### v1

# Dhaka 1229, Bangladesh

Abstract
This paper analyzes the online facility assignment problem in a geometric setting where facilities with unit capacity are positioned at the vertices of a regular $n$-gon. Customers arrive sequentially at uniformly random positions along the edges. They must be assigned immediately to the nearest available facility, with ties broken by coin toss. The sequential nature and unknown future arrivals require a probabilistic analysis of the expected assignment cost. Our main contribution is a recursive characterization of the expected cost: for any occupancy state $S$, the expected remaining cost $V(S)$ equals the average over all edge positions of the immediate assignment cost plus the expected future cost after assignment. We prove that this integral equation can calculate a solution and provide the expected value for small $n$ ($n=3,4,5$). For larger values of $n$ and expected cost, we develop efficient numerical methods, including a discretized dynamic programming approach and Monte Carlo simulation. The work establishes a fundamental probabilistic approach for online assignment in polygonal environments.
1 Introduction
The online facility assignment problem has become increasingly relevant in different areas such as logistics, computer networks, and distributed systems, where requests must be served immediately upon arrival without knowledge of future demand.
1.1 Problem Motivation
The online facility assignment problem in geometric settings captures essential features of real-world systems where service providers with limited capacity must serve customers appearing at random locations. Examples include ride-sharing platforms matching drivers to riders, delivery services allocating packages to vehicles, and sensor networks routing data to processing stations.
1.2 Problem Definition
We consider a regular $n$-gon with facilities at each vertex, each with unit capacity. Customers arrive sequentially according to our assignment system: each customer selects an edge uniformly at random from the $n$ edges, then a position $t\in[0,1]$ uniformly along that edge. Upon arrival, the customer must be assigned to the nearest available facility using the shortest path along the polygon edges. When multiple facilities are equally distant, the customer chooses uniformly at random (Coin Toss). The process terminates after exactly $n$ customers, when all facilities are occupied.
The objective is to minimize the expected total assignment cost, defined as the sum of distances travelled by all customers using a greedy approach. The online nature creates a fundamental scenario that each assignment must be made without future knowledge, yet it must account for how it affects future assignment possibilities.
1.3 Contributions
This paper makes the following key contributions:
  1. i.
    Theoretical Foundation: We derive a recursive integral equation representing the expected cost $V(S)$ for any occupancy state $S\subseteq V$:
    $V(S)=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\left[C_{S,e}(t)+\sum_{m\notin S}^{n}P_{m|S,e}(t)V(S\cup\{m\})\right]dt$
    where $C_{S,e}(t)$ is the immediate cost and $P_{m|S,e}(t)$ are transition probabilities.
  2. ii.
    Algorithmic Solutions: We develop efficient computational methods including exact dynamic programming for moderate $n$ and Monte Carlo simulation for larger instances.
  3. iii.
    Computational Characterization: Through experiments for $n=3$ to $12$, we provide expected costs and analyze scaling behaviour with polygon size.
Our work provides the first exact probabilistic characterization of online facility assignment in polygonal environments, establishing a foundation for analyzing more complex geometric assignment problems.
2 Related Work
The problem of assigning customers to facilities in an online setting sits at the intersection of online algorithms, geometric optimization, and stochastic analysis. The classical Online Facility Location (OFL) problem, where the algorithm must decide whether to open a facility or connect to an existing one, was formalized by Meyerson [11]. He provided a randomized algorithm with a logarithmic competitive ratio, a result that has been refined and generalized for various metric spaces [4]. However, our problem differs from OFL in that facility locations are fixed and capacities are strict. This variant, formally known as the Online Facility Assignment (OFA) problem, has been studied by Ahmed et al. [1], who proved that a greedy strategy yields a linear competitive ratio on line and graph geometries. Recently, Malik et al. [8] extended this to regular polygon geometries, focusing on worst-case bounds.
Our work complements these competitive analyses by addressing the average-case scenario, which shares theoretical roots with the Online Bipartite Matching problem. The seminal work by Karp et al. [7] established that for adversarial arrivals, the optimal competitive ratio is 11/e1-1/e. However, when the assumption of adversarial arrivals is relaxed to random order arrivals (ROA), algorithms often perform significantly better. Goel and Mehta [5] demonstrated that the competitive ratio approaches 1 as the size of the instance grows under random permutations. For weighted variants involving distances, Kalyanasundaram and Pruhs [6] showed that simple greedy algorithms can fail in adversarial settings but perform well stochastically. A comprehensive survey of these online matching techniques and their applications in resource allocation is provided by Mehta [9].
In the specific context of geometric probability, the expected cost of matching random points has a deep theoretical concept. The result discussed by Ajtai, Komlós, and Tusnády [2] says that for $n$ random points in a unit square, the optimal matching cost scales as $\Theta(\sqrt{n\log n})$. However, their work provides asymptotic bounds for large $n$ and does not offer methods for calculating exact costs for finite $n$. Furthermore, as the state space in these problems grows exponentially, the determination of exact analytical solutions becomes computationally intractable. To address this, Monte Carlo methods [10] are frequently employed to approximate high-dimensional integrals and expectations in stochastic systems. Caflisch [3] highlights that for problems where standard numerical integration fails due to dimensionality, Monte Carlo convergence rates remain independent of dimension. Our work synthesizes these approaches: we rely on the recursive structure for exactness in small instances and leverage Monte Carlo integration to estimate the expected costs for larger polygon sizes.
3 Problem Formulation
Before analyzing the expected cost, we need to formalize the problem setting. This mathematical model defines the polygon structure, the arrival distribution, and the distance rules that will be used throughout the paper.
Mathematical Model
We consider a regular $n$-gon with vertex set $V=\{0,1,\ldots,n-1\}$ and edge set $E=\{(i,(i+1)\bmod n):i=0,\ldots,n-1\}$. Each vertex hosts a facility with unit capacity, meaning each facility can serve exactly one customer.
Figure 1: Distance Calculation Method
Customers arrive according to a uniform stochastic process: each customer independently selects an edge $e\in E$ with probability $\frac{1}{n}$, then a position $t\in[0,1]$ uniformly along that edge. Here $t=0$ corresponds to the starting vertex of the edge and $t=1$ to the ending vertex.
The distance between a customer at position $t$ on edge $e=(v_{i},v_{j})$ and a facility at vertex $v$ is measured using the shortest path along edges:
$d(e,t,v)=\min\left(t+\delta(v_{i},v),\ (1-t)+\delta(v_{j},v)\right)$
where $\delta(u,v)=\min(|u-v|,n-|u-v|)$ is the graph distance between vertices $u$ and $v$.
According to the Fig. 1 customer $C_{0}$ randomly arrived in between $e=(2,3)$ edge and also select randomly $t$ position. Assuming customer $C_{0}$ is closer to $F_{3}$ facility so (t<(1t))(t<(1-t)). Here, $F_{0}$ is only a free facility so now distance from facility $F_{2}$ to $F_{0}$ is $\delta(F_{0},F_{2})=2$ (each edge is 1 unit) and $\delta(F_{0},F_{3})=2$. Here,
$d(e,t,0)$ $\implies\min\left(t+\delta(F_{3},F_{0}),\ (1-t)+\delta(F_{2},F_{0})\right)$
$\implies\min\left(t+2,\ (1-t)+2\right)$
$\implies t+2$
Using this process, we calculate the distance.
3.1 Online Assignment Process
The assignment process proceeds sequentially as follows:
  1. i.
    A customer arrives at random edge $e\in E$ and position $t\in[0,1]$
  2. ii.
    The customer is assigned to the nearest available facility using the edge based distance.
  3. iii.
    If multiple facilities are equally distant, the customer chooses uniformly at random among them (Coin Toss).
  4. iv.
    The set of occupied facilities is updated to include the chosen facility.
  5. v.
    The process repeats until all $n$ facilities are occupied.
The objective is to minimize the expected total distance travelled by all customers, where the expectation is taken over the random arrival positions and any tie-breaking randomization.
4 Theoretical Analysis
To analyze the behaviour of the online assignment process, it is essential to express the expected total cost in a recursive form. Therefore, in the following subsection, we derive the expected cost recurrence that links each occupancy state to its future evolution.
4.1 Expected Cost Recurrence
The foundation of our analysis is a recursive equation that characterizes the expected total assignment cost from any given state of facility occupancy. Let $V(S)$ represent the expected total cost starting from state $S\subseteq V$ (where $S$ is the set of occupied facilities) until all facilities become occupied
Theorem 4.1
For any state $S\subset V$ with $|S|<n$, the expected total cost satisfies the recurrence:
$V(S)=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\left[C_{S,e}(t)+\sum_{m=1}^{n}P_{m|S,e}(t)\cdot V(S\cup\{m\})\right]dt$
with the boundary condition $V(V)=0$. Here:
  1. i.
    $C_{S,e}(t)=\min\limits_{m\notin S}d(e,t,m)$ is the immediate cost—the distance to the nearest available facility when a customer arrives at position $t$ on edge $e$.
  2. ii.
    $P_{m|S,e}(t)$ is the probability that facility $m$ is chosen given the current state $S$, arrival edge $e$, and position $t$.
  3. iii.
    The sum $\sum_{m=1}^{n}P_{m|S,e}(t)V(S\cup\{m\})$ represents the expected future cost after the current assignment.
Proof
We derive this recurrence through careful conditioning on the random arrival process:
Step 1: Total Expectation. Since customers arrive uniformly over edges and positions:
$V(S)=\mathbb{E}[\text{Total Cost}\mid S]=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\mathbb{E}[\text{Total Cost}\mid S,e,t]dt$
The factor $\frac{1}{n}$ is for uniform edge selection, and the integral averages over all possible positions along each edge.
Step 2: Cost Decomposition. For a fixed arrival at $(e,t)$, the total cost decomposes into immediate and future components:
$\mathbb{E}[\text{Total Cost}\mid S,e,t]=\underbrace{C_{S,e}(t)}_{\text{immediate cost}}+\underbrace{\mathbb{E}[\text{Future Cost}\mid S,e,t]}_{\text{remaining cost}}$
The immediate cost $C_{S,e}(t)$ is deterministic given $(S,e,t)$, as it’s simply the minimum distance to any available facility.
Step 3: Future Cost Expectation. The future cost depends on which facility gets occupied:
$\mathbb{E}[\text{Future Cost}\mid S,e,t]=\sum_{m=1}^{n}\mathbb{P}(\text{choose }m\mid S,e,t)\cdot\mathbb{E}[\text{Total Cost}\mid S\cup\{m\}]$
By definition,
$\mathbb{P}(\text{choose }m\mid S,e,t)=P_{m|S,e}(t)$ and $\mathbb{E}[\text{Total Cost}\mid S\cup\{m\}]=V(S\cup\{m\})$.
Step 4: Boundary Condition. When all facilities are occupied ($S=V$), no further costs are incurred, so $V(V)=0$. Combining these steps yields the stated recurrence.
4.2 Monte Carlo Interpretation of the Recurrence
Although Theorem 4.1 expresses the expected cost as an exact integral, our implementation evaluates this recurrence using a Monte Carlo estimator. Instead of computing
$V(S)=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\big[C_{S,e}(t)+\sum_{m\notin S}P_{m\mid S,e}(t)\,V(S\cup\{m\})\big]\;dt$
Analytically, the program approximates the expectation
$V(S)=\mathbb{E}\big[C_{S,E}(T)+V(S^{+})\big],$
where $E$ is a uniformly random edge and $T\sim\mathrm{Uniform}[0,1]$ is a random position on that edge.
For each of the $K$ Monte Carlo samples, the algorithm:
  1. i.
    selects a random edge $E$ uniformly from $\{0,\dots,n-1\}$,
  2. ii.
    selects a random position $T\in[0,1]$,
  3. iii.
    computes the immediate cost $C_{S,E}(T)$,
  4. iv.
    determines the nearest available facility using the greedy rule (coin-toss tie breaking),
  5. v.
    obtains the next state $S^{+}=S\cup\{m\}$,
  6. vi.
    forms the sample value
    $Y=C_{S,E}(T)+V(S^{+}),$
    where $V(S^{+})$ has already been computed by dynamic programming.
The estimator for $V(S)$ is the sample average
$\widehat{V}(S)=\frac{1}{K}\sum_{k=1}^{K}Y_{k},$
which is an unbiased Monte Carlo approximation of the recurrence in Theorem 4.1. The DP ordering guarantees that all future-state values $V(S^{+})$ are known before evaluating $V(S)$. Thus, the code implements the theoretical recurrence exactly, with the integral replaced by numerical sampling.
4.3 Structural Properties
Recurrence serves several important mathematical properties that provide intuition about the problem structure.
Lemma 1(Probability Normalization)
For all states $S$, edges $e$, and positions $t$, the transition probabilities satisfy:
$\sum_{m=1}^{n}P_{m|S,e}(t)=1$
Proof
This follows directly from the assignment rules. For any arrival $(e,t)$, the customer must choose exactly one facility. If there’s a unique nearest facility, $P_{m|S,e}(t)=1$ for that facility and 0 for others. If there are $k$ equally distant facilities, each has probability $\frac{1}{k}$, and these probabilities sum to $1$.
4.4 Special Case Analysis
To build intuition, we examine important special cases that reveal the recurrence’s structure.
Corollary 1(Both Endpoints Available)
When a customer arrives on edge $e=(i,j)$ and both endpoints are available ($i,j\notin S$), the recurrence simplifies to:
$\frac{1}{n}\int_{0}^{1}\left[\min(t,1-t)+\frac{V(S\cup\{i\})+V(S\cup\{j\})}{2}\right]dt$
for $t\neq\frac{1}{2}$, with appropriate tie-handling at the midpoint.
Proof
When both endpoints are available and $t<\frac{1}{2}$, facility $i$ is closer and is chosen with probability 1. When $t>\frac{1}{2}$, facility $j$ is chosen. At $t=\frac{1}{2}$, both are equally distant and chosen with probability $\frac{1}{2}$ each. The immediate cost is $\min(t,1-t)$, and the future cost averages the two possible next states.
Corollary 2(One Endpoint Available)
If only one endpoint $i$ of edge $e=(i,j)$ is available ($i\notin S,j\in S$), the contribution becomes:
$\frac{1}{n}\int_{0}^{1}\left[t+V(S\cup\{i\})\right]dt$
since the customer must choose facility $i$ regardless of position, with immediate cost $t$.
5 Computational Results
This section presents a comprehensive experimental validation of our theoretical framework through both Monte Carlo simulation and dynamic programming. We examine the expected assignment costs across various regular polygons, and verify the statistical reliability of our results.
5.1 Worked example: Exact expected cost for a square ($n=4$)
We give a complete analytic computation of the expected total assignment cost for the square ($n=4$), showing the integral and probability calculations that implement the recurrence of Theorem 4.1. Label the vertices $V=\{0,1,2,3\}$ clockwise and edges $E=\{(0,1),(1,2),(2,3),(3,0)\}$. For a random arrival on an edge at position $t\in[0,1]$.
Let $C_{k}$ be the immediate assignment cost of the $k$-th arrival. The total expected cost is $V_{4}=\mathbb{E}[C_{1}]+\mathbb{E}[C_{2}]+\mathbb{E}[C_{3}]+\mathbb{E}[C_{4}]$. We compute each term.
First customer.
When all facilities are free, for any edge the immediate cost is $C(t)=\min(t,1-t)$. Thus
$\mathbb{E}[C_{1}]=\int_{0}^{1}\min(t,1-t)\,dt=2\int_{0}^{1/2}t\,dt=\frac{1}{4}.$
Second customer.
Condition on the first chosen vertex being 0 (all vertices symmetric). Free set $\{1,2,3\}$. For each edge we compute the nearest free-facility distance:
$\begin{array}[]{ll}\text{Edge }(0,1):&C(t)=1-t,\quad\int_{0}^{1}(1-t)\,dt=\tfrac{1}{2},\\[2.0pt] \text{Edge }(1,2):&C(t)=\min(t,1-t),\quad\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(2,3):&C(t)=\min(t,1-t),\quad\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(3,0):&C(t)=t,\quad\int_{0}^{1}t\,dt=\tfrac{1}{2}.\end{array}$
Summing and averaging over edges gives
$\mathbb{E}[C_{2}]=\frac{1}{4}\Big(\tfrac{1}{2}+\tfrac{1}{4}+\tfrac{1}{4}+\tfrac{1}{2}\Big)=\frac{3}{8}.$
Third customer.
After the first customer occupies vertex 0, the free facilities are $\{1,2,3\}$. The second customer arrives on a uniformly random edge with a uniform position $t\in[0,1]$. We compute the probability that each free vertex is the nearest facility by evaluating the measure of the regions on each edge where that facility is closest.
Edge (0,1).
Distances: $d(1)=t$, d(2)=2td(2)=2-t, $d(3)=1+t$. Vertex $1$ is always nearest. Thus this edge contributes probability $1/4$ to vertex $1$.
Edge (1,2).
Distances: $d(1)=t$, d(2)=1td(2)=1-t, $d(3)=1+t$. Vertex $1$ is nearest for $t<1/2$ and vertex $2$ is nearest for $t>1/2$. Each receives an interval of length $1/2$, giving probabilities $1/4\cdot 1/2=1/8$ for both $1$ and $2$.
Edge (2,3).
By symmetry, vertices $2$ and $3$ each receive probability $1/8$.
Edge (3,0).
Distances: $d(3)=t$, d(2)=2td(2)=2-t, $d(1)=1+t$. Vertex $3$ is always nearest. Thus it receives probability $1/4$.
Final probabilities.
Summing the contributions gives
$P(1)=\frac{1}{4}+\frac{1}{8}=\frac{3}{8},\qquad P(2)=\frac{1}{8}+\frac{1}{8}=\frac{1}{4},\qquad P(3)=\frac{1}{4}+\frac{1}{8}=\frac{3}{8}.$
These values are used in the computation of $E[C_{3}]$
  1. 1.
    $S=\{0,1\}$ (free $\{2,3\}$): integrals over edges $\tfrac{5}{2}$, so $\mathbb{E}[C_{3}\mid\{0,1\}]=\tfrac{5}{8}$.
  2. 2.
    $S=\{0,2\}$ (free $\{1,3\}$): integrals sum to $2$, so $\mathbb{E}[C_{3}\mid\{0,2\}]=\tfrac{1}{2}$.
Averaging with the above probabilities gives
$\mathbb{E}[C_{3}]=\frac{3}{8}\cdot\frac{5}{8}+\frac{1}{4}\cdot\frac{1}{2}+\frac{3}{8}\cdot\frac{5}{8}=\frac{19}{32}.$
Fourth (final) customer.
When three facilities are occupied, there is a single free vertex (say 0). The distance from a random arrival to that vertex integrates on the four edges to a total of $4$, so The immediate cost for the final arrival is simply the edgewise distance from a uniformly chosen edge and $t$ to that single free vertex.
Edge integrals when the free vertex is 0:
  1. i.
    Edge $(0,1)$: distance to 0 = $t$. Integral: $\int_{0}^{1}t\,dt=\frac{1}{2}$
  2. ii.
    Edge $(1,2)$: distance to 0 = $t+1$. Integral: $\int_{0}^{1}(t+1)\,dt=\frac{1}{2}+1=\frac{3}{2}$
  3. iii.
    Edge $(2,3)$: distance to 0 = 2t2-t. Integral: $\int_{0}^{1}(2-t)\,dt=2-\frac{1}{2}=\frac{3}{2}$
  4. iv.
    Edge $(3,0)$: distance to 0 = 1t1-t. Integral: $\int_{0}^{1}(1-t)\,dt=\frac{1}{2}$
Sum of integrals: $\frac{1}{2}+\frac{3}{2}+\frac{3}{2}+\frac{1}{2}=4$
Average over 4 edges: $\mathbb{E}[C_{4}]=\frac{4}{4}=1$
Total expected cost.
$V_{4}=\frac{1}{4}+\frac{3}{8}+\frac{19}{32}+1\;=\;\frac{71}{32}=2.21875.$
This exact value matches our DP / Monte Carlo computation (Table 1), validating both the analysis and the implementation.
5.2 Experimental Setup
We implemented both the exact dynamic programming algorithm and Monte Carlo simulation to evaluate the expected costs for regular $n$-gons with $n$ ranging from 3 to 9. All experiments were conducted on a standard computer. The Monte Carlo simulations were run with varying sample sizes ($K=1,000$, $5,000$, and $10,000$) to assess convergence and statistical reliability.
The algorithms were implemented in C++ and compiled with optimization flags enabled. For each polygon size, we computed the expected total assignment cost and derived the per-customer cost by dividing the total by $n$. All distance measurements use the edge-based metric defined in Section 3.
5.3 Expected Cost Analysis
The expected cost varies slightly due to randomness. From the following Table 1
Table 1: Monte Carlo Simulation Results:
Expected Costs for Regular N-gons
$n$ Facilities Expected Total Cost Cost per Customer
K=1, [^0] K=1{,}000 K=5, [^0] K=5{,}000 K=10, [^0] K=10{,}000
3 Triangle 1.406 1.415 1.414 0.471
4 Square 2.220 2.219 2.222 0.556
5 Pentagon 3.137 3.141 3.138 0.628
6 Hexagon 4.175 4.172 4.159 0.693
7 Heptagon 5.270 5.272 5.284 0.755
8 Octagon 6.477 6.498 6.493 0.812
9 Nonagon 7.766 7.777 7.783 0.865
Note: Results show expected total assignment costs from Monte Carlo simulation with $K$ trials. Cost per customer calculated as total cost divided by $n$. All values represent edge-based distance units.
One of the fundamental structural properties of the online facility assignment problem on regular polygons is that the expected total cost increases as the number of facilities increases. By analyzing the values from Table 1, we can say that the expected total cost is proportional to the number of facilities.
5.4 Algorithm Performance
Algorithm 1 Expected Cost Calculation Algorithm using Monte Carlo DP
1:Input: Number of vertices $n$, Number of samples $K$
2:Output: Expected total distance $V[0]$ (from empty state)
3:Initialize value array $V$ of size $2^{n}$ with 0 $\triangleright$ Backward Induction: Iterate from n1n-1 occupied facilities down to 0
4:for $k\leftarrow n-1$ down to 0 do
5:  for each state mask $S$ with $k$ bits set do
6:   Let $F_{free}$ be the set of unoccupied facilities in $S$
7:   $SumCost\leftarrow 0$ $\triangleright$ Monte Carlo Simulation Loop
8:   for $i\leftarrow 1$ to $K$ do
9:     Sample edge $e\sim\text{Uniform}\{0,\dots,n-1\}$
10:     Sample position $t\sim\text{Uniform}[0,1]$ on edge $e$
11:     Compute $d(e,t,f)$ for all $f\in F_{free}$
12:     $d_{\min}\leftarrow\min_{f\in F_{free}}d(e,t,f)$ $\triangleright$ Identify ties for nearest facility
13:     $Ties\leftarrow\{f\in F_{free}\mid d(e,t,f)=d_{\min}\}$ $\triangleright$ Random Tie-Breaking (Coin Toss)
14:     Select facility $f_{next}$ uniformly at random from TiesTies
15:     $S_{next}\leftarrow S\cup\{f_{next}\}$ $\triangleright$ Next state with facility occupied
16:     $CurrentSampleVal\leftarrow d_{\min}+V[S_{next}]$
17:     $SumCost\leftarrow SumCost+CurrentSampleVal$
18:   end for
19:   $V[S]\leftarrow SumCost/K$ $\triangleright$ Store expected value
20:  end for
21:end for
22:return $V[0]$
Here is the detailed complexity analysis for the Monte Carlo Dynamic Programming algorithm.
5.4.1 Time Complexity
The time complexity is determined by the number of states in the DP, the number of Monte Carlo samples per state, and the work done per sample.
  1. i.
    Number of States ($S$): The state is represented by a bitmask of length $n$. The total number of unique states is $2^{n}$.
  2. ii.
    Monte Carlo Samples ($K$): For every single state, the algorithm runs a simulation loop $K$ times.
  3. iii.
    Work Per Sample ($W$): Inside the simulation loop, the algorithm must iterate through all free facilities to calculate distances. In the worst case (at the start), there are $n$ free facilities. Finding the minimum distance and identifying ties requires scanning the list of free facilities once. Therefore, the work per sample is linear with respect to $n$: $W=\mathcal{O}(n)$
The total time complexity is calculated as:
$\text{Time}=\text{States}\times\text{Samples}\times\text{Work per Sample}$
$\text{Time}=\mathcal{O}(2^{n}\cdot K\cdot n)$
5.4.2 Space Complexity
The space complexity is dominated by the storage required for the Dynamic Programming table.
  1. i.
    DP Table ($V$): We store one floating-point value for every possible state mask.
    $\text{Space}_{table}=\mathcal{O}(2^{n})$
  2. ii.
    Auxiliary Storage: The grouping of masks by population count also requires storage proportional to the number of states. Vectors for ties and free facilities take $\mathcal{O}(n)$ space, which is negligible compared to $2^{n}$.
Table 2: Complexity Summary
Metric Complexity Explanation
Time $\mathcal{O}(n\cdot K\cdot 2^{n})$ Process $2^{n}$ states, $K$ simulations each, scanning $n$ facilities.
Space $\mathcal{O}(2^{n})$ Storage required for expected values of all subsets.
5.4.3 Summary and Implications
  1. i.
    Space is dominated by $n$: The dynamic program stores one value for every subset $S\subseteq\{0,\dots,n-1\}$, which yields $2^{n}$ states in total. Thus the memory requirement grows exponentially with $n$, and becomes infeasible beyond $n\approx 20$ on standard hardware ($2^{20}\!\approx 10^{6}$, $2^{30}\!\approx 10^{9}$).
  2. ii.
    Time is dominated by the Monte Carlo parameter $K$: Each state evaluation uses $K$ samples to approximate the integral in the recurrence of Theorem 4.1. Therefore the total runtime is $\Theta(K\,2^{n})$, and increasing $K$ improves accuracy at the cost of a linear increase in computation time.
6 Conclusion
This paper has established a complete theoretical and computational framework for the online facility assignment problem on regular polygons, deriving an exact recursive characterization of expected costs and developing efficient dynamic programming and Monte Carlo solutions that handle the edge-based distance metric and unit capacity constraints. Our analysis reveals fundamental structural properties, including cost monotonicity, and provides the expected total costs for various polygon sizes, offering both practical algorithms for moderate instances and foundational insights. Also identifies some open problems, including non-regular polygons, non-uniform arrivals, higher capacities, and adversarial analysis.
7 Acknowledgements
The authors thank the anonymous reviewers in advance for their valuable feedback. Generative AI tools (ChatGPT and Gemini) were used only for formatting and (Grammarly) for grammatical support, not for technical contributions. The authors declare no conflict of interest.
Appendix 0.A Appendix
0.A.1 Monte Carlo Dynamic Programming Source Code
1#include <bits/stdc++.h>
2#include <random>
3#include <iomanip>
4using namespace std;
5
6// --- Random Number Generator Setup ---
7mt19937 rng(chrono::steady_clock::now().time_since_epoch().count());
8
9double random_double() {
10 return uniform_real_distribution<double>(0.0, 1.0)(rng);
11}
12int random_int(int n) {
13 return uniform_int_distribution<int>(0, n-1)(rng);
14}
15
16// --- Geometry Helpers ---
17double steps_between(int a, int b, int n) {
18 int forward = (b - a + n) % n;
19 int backward = (a - b + n) % n;
20 return min(forward, backward);
21}
22double edgewise_distance(int n, int e, double t, int v) {
23 int v1 = e;
24 int v2 = (e + 1) % n;
25 double dist1 = t + steps_between(v1, v, n);
26 double dist2 = (1.0 - t) + steps_between(v2, v, n);
27 return min(dist1, dist2);
28}
29
30int main() {
31 int n = 9;
32 int NUM_SAMPLES = 5000;
33 cout << fixed << setprecision(3);
34
35 int N = 1 << n;
36 vector<double> V(N, 0.0);
37
38 vector<vector<int>> masks_by_pc(n+1);
39 for (int m=0; m<N; m++)
40 masks_by_pc[__builtin_popcount(m)].push_back(m);
41
42 // === DP BACKWARD INDUCTION ===
43 for (int pc = n-1; pc >= 0; pc--) {
44 for (int mask : masks_by_pc[pc]) {
45 vector<int> freev;
46 for (int v=0; v<n; v++)
47 if (!(mask & (1<<v))) freev.push_back(v);
48
49 double total_cost_accumulator = 0.0;
50
51 for (int k = 0; k < NUM_SAMPLES; k++) {
52 int e = random_int(n);
53 double t = random_double();
54 double min_d = 1e18;
55 vector<int> ties;
56
57 for (int fv : freev) {
58 double d = edgewise_distance(n, e, t, fv);
59 if (d < min_d - 1e-9) {
60 min_d = d;
61 ties.clear();
62 ties.push_back(fv);
63 } else if (abs(d - min_d) < 1e-9) {
64 ties.push_back(fv);
65 }
66 }
67
68 int chosen_facility;
69 if (ties.size() == 1) {
70 chosen_facility = ties[0];
71 } else {
72 int random_index = random_int(ties.size());
73 chosen_facility = ties[random_index];
74 }
75
76 int next_mask = mask | (1 << chosen_facility);
77 double future_val = V[next_mask];
78 double trial_cost = min_d + future_val;
79
80 total_cost_accumulator += trial_cost;
81 }
82 V[mask] = total_cost_accumulator / NUM_SAMPLES;
83 }
84 }
[…]
86 return 0;
87}